OpenStudy (canada907cat):

Determine if the function g(x)={x+1, x<1 and 2 sqrtx, x>1 is differentiable at x=1

OpenStudy (learner):

Basically - do both parts of this piece-wise function have the SAME SLOPE at x=1?

OpenStudy (canada907cat):

ah okay let me do that know, could you check my answer to see if they are right?

OpenStudy (learner):

Sure, go ahead!

OpenStudy (canada907cat):

okay cool thx!

OpenStudy (learner):

I can do an example of a similar problem while you do that (if you want).

OpenStudy (canada907cat):

oh okay yea that be amazing thank you

OpenStudy (learner):

Oh, wait, on thing about your function ... it should say should have one \(\le\) or \(\ge\) sign.

OpenStudy (canada907cat):

oh well it doesn't say so on my assignment

OpenStudy (learner):

So your piece-wise function is \(\displaystyle\color{black}{f(x) =\begin{cases}x+1 & x < 1\\2\displaystyle\sqrt{x} & x > 1\end{cases}}\)

OpenStudy (canada907cat):

yes

OpenStudy (canada907cat):

oh wait its g(x) but I feel that doesn't matter

OpenStudy (learner):

it's weird that there is no point at x=1.

OpenStudy (canada907cat):

sooo what does that mean for answering this problem?

OpenStudy (learner):

You want to know if the function has a slope at \(x=1\), however there is not even a point at \(x=1\).

OpenStudy (canada907cat):

Does that mean the same as differentiable at x=1?

OpenStudy (learner):

In order for a function to be differentiable at \(x=c\) it must be continous at \(x=c\). If \(f\) does not have a point at \(x=c\), then (all the more so) it is not differentiable at \(x=c\).

OpenStudy (canada907cat):

I also know there is a hole at (1, 2) based on the graph soo idk if that a factor in this.

OpenStudy (canada907cat):

so okay okay, but I have to solve it by g(1) or something like that

OpenStudy (learner):

Well, if the piece wise function you are dealing with is: \(\displaystyle\color{blue}{f(x) =\begin{cases}x+1 & x < 1\\2\displaystyle\sqrt{x} & x > 1\end{cases}}\) And not \(\displaystyle\color{black}{f(x) =\begin{cases}x+1 & x < 1\\2\displaystyle\sqrt{x} & x \color{red}{\ge} 1\end{cases}}\) or \(\displaystyle\color{black}{f(x) =\begin{cases}x+1 & x \color{red}{\le} 1\\2\displaystyle\sqrt{x} & x > 1\end{cases}}\)

OpenStudy (learner):

then, you simply don't have a point at x=1. (Your function is not continuous at x=1)

OpenStudy (learner):

And if a function is NOT CONTINUOUS at a point, THEN it is NOT DIFFERENTIABLE at this point.

OpenStudy (canada907cat):

okay so how would I prove that while using the limit definition

OpenStudy (learner):

You don't need to prove that your function is NOT DIFFERENTIABLE at x=1. All you need is the fact that the function is NOT CONTINUOUS at x=1, (that is - there is no point at x=1.)

OpenStudy (learner):

«\(f\) is differentiable at a point x=c» is an equivalent of saying, «\(f\) has a slope (a defined slope) at x=c»

OpenStudy (learner):

If \(f\) does not even contain a point at x=c, then it certainly can not have a (defined) slope at x=c. (Right?)

OpenStudy (learner):

and this is exactly what's going on with your function - it does not contain a point at x=1, and thus it cannot be differentiable at x=1.

OpenStudy (canada907cat):

okay sure it just that the material requires us to prove it using the definition of a limit but I will just write that down in my assignment.

OpenStudy (canada907cat):

Thank you for your help I have a few more if you dont mind helping me with it.

OpenStudy (learner):

More support: Definition: *A differentiable function of one real variable is a function whose derivative exists at each point in its domain.* (and x=1, is apparently not even in the domain of \(f\))

OpenStudy (learner):

Again, unless there was a typo and a function has a \(\le\) or \(\ge\) in one of the restrictions of the piece wise function (as I pointed out). (because that would dictate continuity at x=1, since the parts are equivalent when x=1 .... if there was x=1 in the domain, which there isn't) but, for other than that .... You Welcome! ~ Bye